jhonyy9 (jhonyy9):

Prove this trig identity csc^4-csc^2=cot^4 cot^2 csc = 1/sin csc^4 -csc^2 = cot^4 *cot^2 cot = cos/sin 1/sin^4 - 1/sin^2 = (cos/sin)^4 *(cos/sin)^2 (1 - sin^2) -------------- = (cos^4)/(sin^4) *(cos^2)/(sin^2) sin^4 sin^2 +cos^2 = 1 so from this result 1-sin^2 = cos^2 cos^2 cos^6 ---------- = --------- cross multiply sin^4 sin^6 sin^6 *cos^2 = sin^4 *cos^6 divide both sides by sin^4*cos^2 sin^2 = cos^4 is this far right ?

jhonyy9 (jhonyy9):

@TheSmartOne @ganeshie8 @Kainui @Nnesha

jhonyy9 (jhonyy9):

@zepdrix

OpenStudy (will.h):

@agent0smith

jhonyy9 (jhonyy9):

ty. Will.H

OpenStudy (will.h):

My pleasure

jhonyy9 (jhonyy9):

@Conqueror @zpupster

jhonyy9 (jhonyy9):

@sweetburger

jhonyy9 (jhonyy9):

@.Sam.

jhonyy9 (jhonyy9):

@AloneS

OpenStudy (sshayer):

\[\csc ^4x-\csc ^2x=\csc ^2x \left( \csc ^2x-1 \right)=\csc ^2x \cot ^2x\]

OpenStudy (alivejeremy):

O_O

OpenStudy (sshayer):

i think there is something wrong in the statement.

OpenStudy (sshayer):

\[\csc ^2x \cot ^2x=\left( 1+\cot ^2x \right)\cot ^2x=\cot ^2x+\cot ^4x\]

OpenStudy (kevin):

I think, you don't need to calculate both sides You can just calculating left or right side

jhonyy9 (jhonyy9):

yes Kevin i think it same you - bc. we need proving that the right or left side is equivalent with the other one

OpenStudy (kevin):

I will prove the left: csc^4-csc^2 = 1/sin^4 - 1/sin^2 = 1/sin^4 - sin^2/sin^4 = (1-sin^2)/sin^4

OpenStudy (kevin):

(1-sin^2)/sin^4 = cos^2 / sin^4 = (cos^4 + cos^2sin^2)/ sin^4 = cos^4/sin^4 + (cos^2sin^2)/sin^4 = cot^4 + cot^2

jhonyy9 (jhonyy9):

@Kevin but there in the first line what i posted above there are on the right side cot^4 time cot^2 - so time not plus

OpenStudy (kevin):

what o_O

OpenStudy (kevin):

was your question correct?

jhonyy9 (jhonyy9):

i ve got it on the other math forum and i solved it till this step what you see on the top

OpenStudy (mathmate):

@jhonyy9 You may want to verify your source. See @sshayer's post above. He is the first one to post a correct solution.

jhonyy9 (jhonyy9):

ok. ty.